### v1

# Context-Aware Deep Learning for Robust Channel Extrapolation in Fluid Antenna Systems

Abstract
Fluid antenna systems (FAS) offer remarkable spatial flexibility but face significant challenges in acquiring high-resolution channel state information (CSI), leading to considerable overhead. To address this issue, we propose CANet, a robust deep learning model for channel extrapolation in FAS. CANet combines context-adaptive modeling with a cross-scale attention mechanism and is built on a ConvNeXt v2 backbone to improve extrapolation accuracy for unobserved antenna ports. To further enhance robustness, we introduce a novel spatial amplitude perturbation strategy, inspired by frequency-domain augmentation techniques in image processing. This motivates the incorporation of a Fourier-domain loss function, capturing frequency-domain consistency, alongside a spectral structure consistency loss that reinforces learning stability under perturbations. Our simulation results demonstrate that CANet outperforms benchmark models across a wide range of signal-to-noise ratio (SNR) levels.
Index Terms:
Fluid antenna system (FAS), channel extrapolation, deep learning, amplitude perturbation.
I Introduction
With the advent of sixth-generation (6G) mobile communications, the demand for higher capacity, wider coverage, and greater flexibility in wireless systems is accelerating. Meeting these ambitious requirements necessitates the development of advanced physical-layer technologies. Among them, the fluid antenna system (FAS) [1, 2] has emerged as a promising antenna architecture that introduces new spatial degrees of freedom (DoF) into wireless communication. FAS represents the capability of shape-and-position-reconfigurable antenna technology to empower the physical layer. By leveraging mechanical motion [3], liquid metals [4], reconfigurable pixel-based structures [5] or metamaterials [6], FAS enables dynamic port selection from a massive number of predefined locations. This adaptability facilitates spatial diversity gains in nearly continuous space and enhances the system’s capability to respond intelligently to the wireless channel. In [7], Lu et al. explained FAS from the viewpoint of electromagnetics.
Since the introduction of FAS by Wong et al. in [8, 9], there have been lots of efforts in understanding the achievable performance of FAS [10, 11]. The use of FAS has also been considered for secrecy communications [12], reconfigurable intelligent surface (RIS) systems [13], index modulation [14, 15], wideband communication [16], and integrated sensing and communication (ISAC) [17]. Capacity maximization for FAS channels has also been studied using statistical channel state information (CSI) [18] and instantaneous CSI [19]. The diversity-multiplexing trade-off for multiple-input multiple-output (MIMO) FAS channels was analyzed in [20].
Nonetheless, the promising performance of FAS relies on the availability of CSI and the high-resolution nature of FAS also leads to substantial overhead in acquiring CSI, including pilot overhead, feedback latency, and energy consumption. In recent years, several studies explored the integration of physical modeling and data-driven approaches for CSI prediction [21, 22]. For instance, [23] proposed an inference algorithm to enable high-accuracy extrapolation in a massive FAS using only a small number of observed ports. Additionally, in [24], an extrapolation scheme based on masked language models, which reconstructs the full channel from a limited number of observed port CSI, was proposed. Going one step further, the approach in [25] integrates the diffusion mechanism of graph neural networks with an asymmetric masked autoencoder architecture, significantly improving both extrapolation accuracy and flexibility in high-resolution FAS settings. Prior to these, there were attempts that exploited channel sparsity [26, 27] and used Bayesian approaches [28, 29]. CSI prediction for multiuser FAS using deep learning was also reported in [30]. Overall, however, reliable CSI acquisition seems only possible at high signal-to-noise ratio (SNR) and it is important to study enhancement techniques that work well in noisy situations.
Motivated by the above, in this correspondence, we propose CANet, a novel deep learning framework designed for channel extrapolation in FAS. CANet adaptively infers the CSI of the unobserved ports with varying positions within a FAS, and maintains robust performance across different SNR scenarios. The architecture of CANet incorporates a context adaptive block (CAB), a cross-scale contextual attention (CSCA) mechanism, and a ConvNeXt v2 backbone enhanced with dropout regularization. This combination empowers CANet to process unobserved port positions both locally and globally, significantly improving the extraction of CSI features. To enhance model robustness further, we also introduce a spatial amplitude perturbation strategy. Besides, we refine the loss function by incorporating a fast Fourier transform (FFT) loss, which collaborates with a spectral structure consistency loss to reinforce model stability. Numerical results indicate that CANet surpasses benchmark models, demonstrating superior channel extrapolation robustness across varying noise levels.
II Channel Extrapolation for FAS
Consider a downlink system where a base station (BS) with $M_{t}$ fixed-position antennas (FPAs) communicates with $U_{t}$ user terminals on the same time-frequency resource unit. Each user terminal employs a 2D FAS with $N_{s}=N_{x}\times N_{y}$ switchable positions (or ports), which are evenly distributed over a planar surface of physical dimensions $W_{s}=W_{x}\times W_{y}$. In a 3D rich scattering environment, the spatial correlation between port $(n_{x}^{s},n_{y}^{s})$ and port $(\tilde{n}_{x}^{s},\tilde{n}_{y}^{s})$ is given by
$\begin{array}[]{l}J_{\left(n_{x},n_{y}\right),\left(\tilde{n}_{x},\tilde{n}_{y}\right)}\\ \quad=j_{0}\left(\frac{2\pi}{\lambda}\sqrt{\left(\frac{\left|n_{x}-\tilde{n}_{x}\right|}{N_{x}-1}W_{x}\right)^{2}+\left(\frac{\left|n_{y}-\tilde{n}_{y}\right|}{N_{y}-1}W_{y}\right)^{2}}\right),\end{array}$ (1)
where $j_{0}(\cdot)$ denotes the spherical Bessel function of the first kind, which can be approximated by $\mathrm{sinc}(x)={\sin x}/{x}$, and $\lambda$ represents the carrier wavelength. The downlink channel for a typical user can then be modeled as
$\mathbf{g}=\delta\mathbf{U}_{s}\sqrt{\mathbf{\Lambda}_{s}}\mathbf{G}+\mathbf{n},$ (2)
where $\mathbf{U}_{s}\in\mathbb{C}^{N_{s}\times N_{s}}$ and $\mathbf{\Lambda}_{s}\in\mathbb{C}^{N_{s}\times N_{s}}$ are defined in accordance with [20], representing the eigenvector matrix and eigenvalue diagonal matrix obtained from the eigen-decomposition of the spatial correlation matrix $\mathbf{J}_{s}=\mathbf{U}_{s}\mathbf{\Lambda}_{s}\mathbf{U}_{s}^{H}$ , whose entries are given by (1). The matrix $\mathbf{G}\in\mathbb{C}^{N_{s}\times M_{t}}$consists of independent complex Gaussian random variables with zero mean and variance of $1/2$. The parameter $\delta$ characterizes the path loss, and ${\bf n}$ is the additive white Gaussian noise.
III Proposed Approach
Here, we propose CANet, a noise-robust architecture for the channel extrapolation task. We design a CAB to locally fuse the features of known CSI with weighted integration, while constraining the propagation of features from unknown CSI. To enhance global modeling capability, we introduce a CSCA module to facilitate feature completion for large unobserved port regions. Moreover, we adopt ConvNeXt v2 with dropout regularization to improve feature extraction and generalization after feature completion [31]. A spatial amplitude perturbation module is also incorporated to enhance the model’s adaptability and robustness under various noise conditions.
III-A Masking Strategy
We denote the perfect CSI as $\mathbf{g}\in\mathbb{C}^{N_{s}\times M_{t}}$, and construct a tensor $\mathbf{u}\in\mathbb{R}^{2M_{t}\times N_{y}\times N_{x}}$ to represent the concatenated real and imaginary components of the CSI across a 2D fluid antenna grid of size $N_{x}\times N_{y}$. To simulate partial observations, a subset of antenna ports is randomly masked along the spatial grid, with a masking ratio typically ranging from $80\%$ to $95\%$. The resulting masked tensor is denoted as $\tilde{\mathbf{u}}\in\mathbb{R}^{2M_{t}\times N_{y}\times N_{x}}$, in which the unobserved positions are replaced with a distinct negative value (e.g., $-10$) to make them distinguishable for the network. Additionally, we define a binary indicator matrix $\mathbf{E}_{\text{flag}}\in\{0,1\}^{N_{y}\times N_{x}}$, where a value of $1$ indicates an observed port and 00 is a masked (unobserved) port. No prior knowledge of the channel is used to select ports for extrapolation. Instead, a random masking strategy is employed to ensure flexibility and enhance the generalization capability of the model.
Figure 1: The proposed CANet for CSI extrapolation in FAS channels, where $\oplus$, $\otimes$ and $\odot$ denote fuse, dot product and element-wise product, respectively.
III-B Model Architecture
The overall architecture of CANet consists of four training modules. Next, we summarize the role of each module.
III-B1 CAB
For channel extrapolation tasks with strong spatial correlations, using convolutional neural network is a natural choice for feature extraction. However, conventional convolution is susceptible to interference from unobserved ports within the receptive field, which can degrade the quality of feature representations. To overcome this issue, we propose a dual-branch architecture that separately models and predicts the CSI of observed and unobserved ports, thereby enhancing both the modeling accuracy and the robustness of extrapolation.
Specifically, we first compute a soft mask (confidence map) over the entire feature map using convolution. This soft mask reflects the reliability of each feature. [^1] Features affected by masked ports are assigned lower confidence, while those from observed regions receive higher confidence. Then we perform a weighted fusion of the outputs from the convolutional and attention branches to further refine feature extraction within the receptive field. For each pixel $x_{o}\in\mathbb{R}^{C}$ on the feature map $F\in\mathbb{R}^{C\times H\times W}$, its corresponding soft mask value $m$ is computed by the following convolutional operation:
$m=\sigma\left(\sum_{i\in\mathcal{B}_{o}}w_{i}^{m}x_{i}\right),$ (3)
where $\sigma$ denotes the sigmoid activation function, $x_{i}$ represents the input features within the receptive field, $w_{i}$ denotes learnable convolutional filters, and $\mathcal{B}_{o}$ represents the receptive field range corresponding to filter $w^{m}$. The magnitude of $m$ signifies the confidence level of the feature at that location. Conditioned on the learned soft mask $m$, we then introduce a self-attention mechanism [32] to derive a more contextually relevant feature representation $x^{a}$, given by
$x^{a}=\sum_{i\in\mathcal{B}_{o}}\left\langle mw^{q}x_{o},mw^{k}x_{i}\right\rangle mx_{i}w^{v}x_{i},$ (4)
[…]
$m\in\sum_{i\in\mathcal{B}_{o}}w_{i}^{c}x_{i}.$ (5)
Finally, we perform a weighted fusion of the convolutional branch output $x_{c}$ and the attention branch output $x_{a}$ based on the soft mask $m$, yielding the output
$o=m\cdot x^{c}+(1-m)\cdot x^{a}.$ (6)
III-B2 CSCA
In the CSCA module, we first reduce the input feature channels to half of the original before feeding them into two self-attention modules. One branch directly undergoes self-attention at the same scale, while the other duplicates the feature map and down-samples it to half of the original resolution. The downsampled feature map is then processed along with the original-scale features within the self-attention mechanism. At each self-attention stage, patches are extracted from the input features, and the cosine similarity between patches in the missing region and those in the external region is computed to establish cross-region feature correlations. Formally, let $q_{i}$ denote the $i$-th patch from the observed region $\psi^{l}$, and $k_{j}$ denote the $j$-th patch from the masked region $\phi^{l}$. The cosine similarity is computed as
$s_{i,j}=\left\langle\frac{q_{i}}{\|q_{i}\|},\frac{k_{j}}{\|k_{j}\|}\right\rangle.$ (7)
The attention weights are then computed by
$a_{i,j}=\frac{\exp(s_{i,j})}{\sum_{j}\exp(s_{i,j})}.$ (8)
The unobserved port patches are then reconstructed based on the computed attention scores, given by
$\hat{q}=\sum_{j}a_{i,j}v_{j},$ (9)
where $v_{j}$ represents the $j$-th patch extracted from features outside unobserved regions. Finally we concatenate the results of the two branches as the final result.
TABLE I: Architecture of the proposed CANet
Layer Operator Input Size Output Size Output Channel Kernel Size Stride Nonlinearity
1 CAB $32~{}\times~{}16$ $32~{}\times~{}16$ $64$ $5$ $1$ LeakyReLU
2 CAB $32~{}\times~{}16$ $32~{}\times~{}16$ $128$ $3$ $2$ LeakyReLU
3 CAB $32~{}\times~{}6$ $16~{}\times~{}8$ $128$ $3$ $1$ LeakyReLU
4 CAB $16~{}\times~{}8$ $16~{}\times~{}8$ $256$ $3$ $1$ LeakyReLU
5 CAB $16~{}\times~{}8$ $16~{}\times~{}8$ $256$ $3$ $1$ LeakyReLU
6 CSCA $16~{}\times~{}8$ $16~{}\times~{}8$ $256$ Null Null Null
7–16 ConvNeXt v2 $16~{}\times~{}8$ $16~{}\times~{}8$ $256$ $7$ $3$ GeLU
17 BilinerUp $16~{}\times~{}8$ $32~{}\times~{}16$ $256$ $3$ $1$ Null
18 CAB $32~{}\times~{}16$ $32~{}\times~{}16$ $128$ $3$ $1$ LeakyReLU
19 CAB $32~{}\times~{}16$ $32~{}\times~{}16$ $128$ $3$ $1$ LeakyReLU
20 CAB $32~{}\times~{}16$ $32~{}\times~{}16$ $64$ $3$ $1$ LeakyReLU
21 CAB $32~{}\times~{}16$ $32~{}\times~{}16$ $64$ $3$ $1$ LeakyReLU
22 Conv $32~{}\times~{}16$ $32~{}\times~{}16$ $16$ $7$ $1$ Null
III-B3 ConvNeXt v2 with dropout regularization
After CAB and CSCA, we adopt ConvNeXt V2 [31] with dropout regularization as the backbone network to further enhance feature extraction and improve generalization. ConvNeXt has been widely used for its strong feature representation capabilities, and ConvNeXt V2 builds upon this by introducing the global response normalization (GRN) module, which enables more expressive feature representations. In our design, we employ the ConvNeXt V2 block for channel extrapolation. By randomly dropping a subset of neurons during training, dropout alters the effective network structure in each training epoch, leading to a degree of statistical independence across different training instances, thereby reducing prediction variance and enhancing generalization under diverse noise conditions.
III-B4 Spatial amplitude perturbation
To enhance the model’s robustness, we propose a spatial amplitude perturbation strategy at the input stage, inspired by frequency perturbation techniques in the image domain [33, 34]. Specifically, the input CSI features are first zero-mean normalized to remove the DC component, then transformed into the frequency domain via FFT. The amplitude spectrum is then perturbed in a stochastic manner based on a predefined probability, effectively simulating potential frequency interference and noise. This approach improves the model’s adaptability to unseen distributions:
$\tilde{\mathbf{z}}=\tilde{\mathbf{u}}-\frac{1}{HW}\sum_{h=1}^{H}\sum_{w=1}^{W}\tilde{\mathbf{u}}_{h,w}.$ (10)
$\hat{\mathbf{z}}=\mathcal{F}^{-1}\left(\mathcal{F}(\tilde{\mathbf{z}})\cdot\left(1+\bm{\gamma}\cdot\bm{\xi}\cdot\bm{\delta}\right)\right)+\bar{\mathbf{z}},$ (11)
where $\tilde{\mathbf{z}}$ is the spatially zero-mean feature, $\hat{\mathbf{z}}$ is the final perturbed feature after reconstruction, $\mathcal{F}(\cdot)$ and $\mathcal{F}^{-1}(\cdot)$ denote the 2D discrete Fourier transform and its inverse. The perturbation mask $\bm{\delta}\in\{0,1\}^{H\times W}$ above is sampled from a Bernoulli distribution with probability $\mu$, and the amplitude coefficients $\bm{\xi}\in[0,1]^{H\times W}$ are sampled from a normal distribution $\mathcal{N}(0,1)$. The hyperparameters $\mu$ and $\gamma$ control the perturbation probability and maximum strength, respectively. Only the amplitude of each frequency component is perturbed, while the phase is fixed to maintain structural consistency.
III-B5 Loss Function
Since the CSI tensor is represented by concatenating the real and imaginary parts, we reconstruct it into a complex-valued tensor before computing the frequency-domain loss to enable the 2D Fourier transform. We denote the original CSI tensor as $\mathbf{{U}_{c}}\in\mathbb{R}^{M_{t}\times N_{y}\times N_{x}}$, and $\hat{\mathbf{{U}_{c}}}\in\mathbb{R}^{M_{t}\times N_{y}\times N_{x}}$ denote the predicted CSI recovered from the masked input $\tilde{\mathbf{{U}_{c}}}$. Let $\Omega\subset\{1,\dots,N_{y}\}\times\{1,\dots,N_{x}\}$ denote the set of unobserved spatial locations, with $|\Omega|=N_{a}$. The total loss is then defined as
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{MSE}}+\beta\mathcal{L}_{\mathrm{FFT}},$ (12)
where $\mathcal{L}_{\text{MSE}}$ and $\mathcal{L}_{\mathrm{FFT}}$ are the mean-square-error (MSE) loss and the FFT loss attributed to spatial amplitude perturbation, respectively, calculated as
$\mathcal{L}_{\text{MSE}}$ $=\mathbb{E}\left[\frac{1}{|\Omega|}\sum_{(y,x)\in\Omega}\left\|\hat{\mathbf{{U}_{c}}}[:,y,x]-\mathbf{{U}_{c}}[:,y,x]\right\|_{2}^{2}\right],$ (13)
$\mathcal{L}_{\mathrm{FFT}}$ $=\mathbb{E}\left[\left\|\left|\mathcal{F}(\hat{\mathbf{{U}_{c}}})\right|-\left|\mathcal{F}(\mathbf{{U}_{c}})\right|\right\|_{2}^{2}\right],$ (14)
in which $\beta$ balances the contribution of the two losses. The FFT loss enforces global consistency in the spatial frequency domain, which helps preserve large-scale channel structure.
IV Numerical Results
In the simulations, the FAS channel involves $U_{t}$ users. Each user has a planar FAS of size $W_{s}\,~{}\text{cm}^{2}$ and $N_{s}$ antenna ports. We assume $m_{t}=U_{t}$, and the BS is equipped with $M_{t}$ FPAs. The carrier frequency is $3.4~{}{\rm GHz}$. Detailed simulation parameters are summarized in Table I. The dataset used for analysis contains a total of $310,000$ samples, divided into $240,000$ training samples, $60,000$ validation samples, and $10,000$ test samples for model training and performance evaluation. The total number of ports at the FAS side is set to $N_{s}=512$.
The performance is evaluated through the normalized MSE (NMSE), defined as
$\text{NMSE}(\hat{\mathbf{g}},\mathbf{g})=\frac{\sum_{n=1}^{N_{\text{test}}}\left\|\mathbf{g}_{n}-\hat{\mathbf{g}}_{n}\right\|^{2}}{\sum_{n=1}^{N_{\text{test}}}\left\|\mathbf{g}_{n}\right\|^{2}},$ (15)
where $N_{\text{test}}$ denotes the number of test samples. The predicted CSI matrix $\hat{\mathbf{g}}$ is obtained by rearranging the extrapolated tensor $\hat{\mathbf{U}}_{c}$ into the standard CSI format. The batch size used in training is set to $32$. The initial learning rate is $0.0005$. For optimization, the AdamW optimizer is employed with the parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, and the weight decay set to $0.0001$. We set the maximum strength $\gamma$ to $0.5$, $\mu$ to $0.05$, and the dropout rate to $0.5$, $\beta$ to $0.02$.
(a) $W_{s}=2~{}{\rm cm}\times 4~{}{\rm cm}$
(b) $W_{s}=8~{}{\rm cm}\times 16~{}{\rm cm}$
Figure 2: NMSE versus the number of observed ports with different FAS sizes.
Figs. 2(a) and 2(b) illustrate the NMSE curves of different models under various SNR conditions, as a function of the observed port ratio, for FAS sizes of $2~{}{\rm cm}\times 4~{}{\rm cm}$ and $8~{}{\rm cm}\times 16~{}{\rm cm}$, respectively. We reproduce the AGMAE in [25] as a baseline for comparison. The simplified version of CANet without spatial amplitude perturbation and spectral consistency loss is denoted as ‘CANet-B’, while the full CANet model includes all the proposed components. In the figures, curves with the same line style represent the same SNR setting.
The results demonstrate that as the ratio of observed ports increases, the NMSE of all the models decreases considerably, showing that richer observed information leads to better reconstruction performance. Under extremely low observation rates (e.g., $5\%$), AGMAE slightly outperforms CANet-B, suggesting that its structure has certain advantages under highly sparse inputs. But in most observation rates, CANet-B consistently surpasses AGMAE. Furthermore, with the incorporation of amplitude spectrum perturbation and spectral consistency loss, the overall performance of CANet is further improved. For example, at an observation rate of $20\%$ with the $2~{}{\rm cm}\times 4~{}{\rm cm}$ FAS, CANet achieves up to a $3~{}{\rm dB}$ NMSE gain over CANet-B, demonstrating stronger noise robustness and generalization ability. Fig. 2(b) shows a similar trend under a larger FAS size, validating the adaptability and effectiveness of the proposed method across different physical FAS configurations.
V Conclusions
In this correspondence, we proposed CANet, a robust CSI extrapolation model that excels in processing unobserved port positions in FAS. CANet integrates a CAB, CSCA, and a ConvNeXt v2 backbone. This combination allows for adaptive CSI feature extraction at both local and global levels. A spatial amplitude perturbation strategy was also proposed to enhance the loss function with an FFT loss. Our simulation results indicated that CANet consistently outperforms benchmark models, demonstrating improved accuracy and remarkable robustness in channel extrapolation under various noise conditions.

### v2

@@
 in [8, 9], there have been lots of efforts in understanding the achievable performance of FAS [10, 11].
-The use of FAS has also been considered for secrecy communications [12], reconfigurable intelligent surface (RIS) systems [13], index modulation [14, 15], wideband communication [16], and integrated sensing and communication (ISAC) [17].
+The use of FAS has also been considered for reconfigurable intelligent surface (RIS) systems [12], index modulation [13, 14], wideband communication [15], and integrated sensing and communication (ISAC) [16].
 Capacity maximization for FAS channels has also been studied using statistical channel state information (CSI) [17] and instantaneous CSI [18].
 The diversity-multiplexing trade-off for multiple-input multiple-output (MIMO) FAS channels was analyzed in [19].
-Nonetheless, the promising performance of FAS relies on the availability of CSI and the high-resolution nature of FAS also leads to substantial overhead in acquiring CSI, including pilot overhead, feedback latency, and energy consumption.
+Nonetheless, the promising performance of FAS relies on the availability of CSI and the high-resolution nature of FAS also leads to substantial overhead in acquiring CSI, including pilot overhead, feedback latency, and energy consumption[20].
 In recent years, several studies explored the integration of physical modeling and data-driven approaches for CSI prediction [21, 22].
@@
 Going one step further, the approach in [25] integrates the diffusion mechanism of graph neural networks with an asymmetric masked autoencoder architecture, significantly improving both extrapolation accuracy and flexibility in high-resolution FAS settings.
-Prior to these, there were attempts that exploited channel sparsity [26, 27] and used Bayesian approaches [28, 29].
+Prior to these, there were attempts that exploited channel sparsity [26] and used Bayesian approaches [27].
 CSI prediction for multiuser FAS using deep learning was also reported in [28].
@@
 2 CAB $32~{}\times~{}16$ $32~{}\times~{}16$ $128$ $3$ $2$ LeakyReLU
-3 CAB $32~{}\times~{}6$ $16~{}\times~{}8$ $128$ $3$ $1$ LeakyReLU
+3 CAB $32~{}\times~{}16$ $16~{}\times~{}8$ $128$ $3$ $1$ LeakyReLU
 4 CAB $16~{}\times~{}8$ $16~{}\times~{}8$ $256$ $3$ $1$ LeakyReLU

### v4

@@
 Fluid antenna systems (FAS) offer remarkable spatial flexibility but face significant challenges in acquiring high-resolution channel state information (CSI), leading to considerable overhead.
-To address this issue, we propose CANet, a robust deep learning model for channel extrapolation in FAS.
+To address this issue, leveraging the inherent spatial continuity of electromagnetic fields, we propose CANet, a robust deep learning model for channel extrapolation in FAS.
 CANet combines context-adaptive modeling with a cross-scale attention mechanism and is built on a ConvNeXt v2 backbone to improve extrapolation accuracy for unobserved antenna ports.
@@
 This motivates the incorporation of a Fourier-domain loss function, capturing frequency-domain consistency, alongside a spectral structure consistency loss that reinforces learning stability under perturbations.
-Our simulation results demonstrate that CANet outperforms benchmark models across a wide range of signal-to-noise ratio (SNR) levels.
+Our simulation results demonstrate that CANet not only outperforms benchmark models in terms of estimation accuracy across a wide range of signal-to-noise ratio (SNR) levels and carrier frequencies, but also significantly reduces the outage probability with moderate computational complexity.
 Index Terms:
@@
 I Introduction
-With the advent of sixth-generation (6G) mobile communications, the demand for higher capacity, wider coverage, and greater flexibility in wireless systems is accelerating.
+With the advent of sixth-generation (6G) mobile communications, the demand for higher capacity, wider coverage, improved positioning and greater flexibility in wireless systems is accelerating[34, 7, 6, 3].
 Meeting these ambitious requirements necessitates the development of advanced physical-layer technologies.
-Among them, the fluid antenna system (FAS) [1, 2] has emerged as a promising antenna architecture that introduces new spatial degrees of freedom (DoF) into wireless communication.
+Among them, the fluid antenna system (FAS) [30, 20, 9] has emerged as a promising antenna architecture that introduces new spatial degrees of freedom (DoF) into wireless communication.
 FAS represents the capability of shape-and-position-reconfigurable antenna technology to empower the physical layer.
-By leveraging mechanical motion [3], liquid metals [4], reconfigurable pixel-based structures [5] or metamaterials [6], FAS enables dynamic port selection from a massive number of predefined locations.
+By leveraging mechanical motion [2], liquid metals [25], reconfigurable pixel-based structures [38] or metamaterials [17, 19], FAS enables dynamic port selection from a massive number of predefined locations.
 This adaptability facilitates spatial diversity gains in nearly continuous space and enhances the system’s capability to respond intelligently to the wireless channel.
-In [7], Lu et al.
-explained FAS from the viewpoint of electromagnetics.
 Since the introduction of FAS by Wong et al.
-in [8, 9], there have been lots of efforts in understanding the achievable performance of FAS [10, 11].
-The use of FAS has also been considered for reconfigurable intelligent surface (RIS) systems [12], index modulation [13, 14], wideband communication [15], and integrated sensing and communication (ISAC) [16].
-Capacity maximization for FAS channels has also been studied using statistical channel state information (CSI) [17] and instantaneous CSI [18].
-The diversity-multiplexing trade-off for multiple-input multiple-output (MIMO) FAS channels was analyzed in [19].
-Nonetheless, the promising performance of FAS relies on the availability of CSI and the high-resolution nature of FAS also leads to substantial overhead in acquiring CSI, including pilot overhead, feedback latency, and energy consumption[20].
-In recent years, several studies explored the integration of physical modeling and data-driven approaches for CSI prediction [21, 22].
+in [28, 29], there have been lots of efforts in understanding the achievable performance of FAS [22, 24], showing potential for reconfigurable intelligent surface (RIS) systems [15], index modulation [40, 41], wideband communication [12], and integrated sensing and communication (ISAC) [26].
+Nonetheless, the promising performance of FAS relies on the availability of CSI [10, 36, 4, 23], and the high-resolution nature of FAS also leads to substantial overhead in acquiring CSI, including pilot overhead, feedback latency, and energy consumption[13, 8].
+In recent years, several studies explored the integration of physical modeling and data-driven approaches for CSI prediction [5, 11, 14].
 For instance, [16] proposed an inference algorithm to enable high-accuracy extrapolation in a massive FAS using only a small number of observed ports.
@@
 CSI prediction for multiuser FAS using deep learning was also reported in [27].
-Overall, however, reliable CSI acquisition seems only possible at high signal-to-noise ratio (SNR) and it is important to study enhancement techniques that work well in noisy situations.
-Motivated by the above, in this correspondence, we propose CANet, a novel deep learning framework designed for channel extrapolation in FAS.
-CANet adaptively infers the CSI of the unobserved ports with varying positions within a FAS, and maintains robust performance across different SNR scenarios.
-The architecture of CANet incorporates a context adaptive block (CAB), a cross-scale contextual attention (CSCA) mechanism, and a ConvNeXt v2 backbone enhanced with dropout regularization.
-This combination empowers CANet to process unobserved port positions both locally and globally, significantly improving the extraction of CSI features.
-To enhance model robustness further, we also introduce a spatial amplitude perturbation strategy.
-Besides, we refine the loss function by incorporating a fast Fourier transform (FFT) loss, which collaborates with a spectral structure consistency loss to reinforce model stability.
-Numerical results indicate that CANet surpasses benchmark models, demonstrating superior channel extrapolation robustness across varying noise levels.
+However, the existing research shows limitations in reliable CSI acquisition for FAS.
+Due to the extremely dense deployment of candidate ports, full-port CSI acquisition via sequential switching incurs prohibitive latency and pilot overhead, making extrapolation from a highly limited number of observed ports inevitable.
+The core challenge lies in reconstructing high-resolution spatial correlation from sparse and noise-contaminated observations, under which conventional learning-based models are prone to noise overfitting and poor generalization to unobserved ports.
+Motivated by these challenges, we propose CANet, a deep learning framework specifically designed for robust channel extrapolation in FAS.
+CANet adopts a synergistic local-to-global architecture with multi-domain regularization, integrating a context-adaptive block (CAB) for reliable local feature fusion, a cross-scale contextual attention (CSCA) mechanism for global feature completion, and a ConvNeXt v2 backbone.
+In addition, a spatial amplitude perturbation strategy and a Fast Fourier Transform (FFT)-based loss are introduced to enhance robustness against noise by enforcing frequency-domain consistency.
 II Channel Extrapolation for FAS
-Consider a downlink system where a base station (BS) with $M_{t}$ fixed-position antennas (FPAs) communicates with $U_{t}$ user terminals on the same time-frequency resource unit.
-Each user terminal employs a 2D FAS with $N_{s}=N_{x}\times N_{y}$ switchable positions (or ports), which are evenly distributed over a planar surface of physical dimensions $W_{s}=W_{x}\times W_{y}$.
+Consider a slow fluid antenna multiple access (FAMA) downlink system[27] where a base station (BS) with $M_{t}$ fixed-position antennas (FPAs) communicates with $U_{t}$ user terminals on the same time-frequency resource unit.
+Each user terminal employs a 2D FAS with $N_{s}=N_{x}\times N_{y}$ switchable positions (or ports), which are evenly distributed over a planar surface of physical dimensions $W_{s}=W_{x}\times W_{y}$, where each user activates only one port in each coherence block.
 In a 3D rich scattering environment, the spatial correlation between port $(n_{x}^{s},n_{y}^{s})$ and port $(\tilde{n}_{x}^{s},\tilde{n}_{y}^{s})$ is given by
-$\begin{array}[]{l}J_{\left(n_{x},n_{y}\right),\left(\tilde{n}_{x},\tilde{n}_{y}\right)}\\ \quad=j_{0}\left(\frac{2\pi}{\lambda}\sqrt{\left(\frac{\left|n_{x}-\tilde{n}_{x}\right|}{N_{x}-1}W_{x}\right)^{2}+\left(\frac{\left|n_{y}-\tilde{n}_{y}\right|}{N_{y}-1}W_{y}\right)^{2}}\right),\end{array}$ (1)
+$\begin{array}[]{l}J_{\left(n_{x},n_{y}\right),\left(\tilde{n}_{x},\tilde{n}_{y}\right)}\\ \qquad=j_{0}\left(\frac{2\pi}{\lambda}\sqrt{\left(\frac{\left|n_{x}-\tilde{n}_{x}\right|}{N_{x}-1}W_{x}\right)^{2}+\left(\frac{\left|n_{y}-\tilde{n}_{y}\right|}{N_{y}-1}W_{y}\right)^{2}}\right),\end{array}$ (1)
 where $j_{0}(\cdot)$ denotes the spherical Bessel function of the first kind, which can be approximated by $\mathrm{sinc}(x)={\sin x}/{x}$, and $\lambda$ represents the carrier wavelength.
-The downlink channel for a typical user can then be modeled as
-$\mathbf{g}=\delta\mathbf{U}_{s}\sqrt{\mathbf{\Lambda}_{s}}\mathbf{G}+\mathbf{n},$ (2)
-where $\mathbf{U}_{s}\in\mathbb{C}^{N_{s}\times N_{s}}$ and $\mathbf{\Lambda}_{s}\in\mathbb{C}^{N_{s}\times N_{s}}$ are defined in accordance with [19], representing the eigenvector matrix and eigenvalue diagonal matrix obtained from the eigen-decomposition of the spatial correlation matrix $\mathbf{J}_{s}=\mathbf{U}_{s}\mathbf{\Lambda}_{s}\mathbf{U}_{s}^{H}$ , whose entries are given by (1).
-The matrix $\mathbf{G}\in\mathbb{C}^{N_{s}\times M_{t}}$consists of independent complex Gaussian random variables with zero mean and variance of $1/2$.
-The parameter $\delta$ characterizes the path loss, and ${\bf n}$ is the additive white Gaussian noise.
-III Proposed Approach
+The same spatial correlation matrix $\mathbf{J}_{s}=\mathbf{U}_{s}\mathbf{\Lambda}_{s}\mathbf{U}_{s}^{H}$ is shared by all users, where $\mathbf{U}_{s}\in\mathbb{C}^{N_{s}\times N_{s}}$ and $\mathbf{\Lambda}_{s}\in\mathbb{C}^{N_{s}\times N_{s}}$ denote the eigenvector matrix and the diagonal eigenvalue matrix, respectively.
+For a given user, the downlink channel over all FAS ports and all FPAs can be modeled as
+$\mathbf{g}=\delta\,\mathbf{U}_{s}\sqrt{\mathbf{\Lambda}_{s}}\mathbf{G},$ (2)
+where $\mathbf{g}\in\mathbb{C}^{N_{s}\times M_{t}}$ denotes the small-scale fading channel matrix, $\mathbf{G}\in\mathbb{C}^{N_{s}\times M_{t}}$ consists of independent complex Gaussian random variables with zero mean and variance $1/2$ per real dimension, and $\delta$ characterizes the large-scale path loss.
+For the $u$-th user terminal, $u\in\{1,\dots,U_{t}\}$, its channel realization over all FAS ports and BS antennas is denoted by $\mathbf{g}_{u}\in\mathbb{C}^{N_{s}\times M_{t}}$, which follows the same statistical model as in (2).
+As demonstrated in [1], if the position-to-channel mapping is bijective in a static FAS environment (with fixed geometry, port locations, and scattering conditions), then a channel-to-channel mapping also exists.
+In other words, once the environment is fixed, each FAS port position almost surely corresponds to a unique downlink channel vector, and vice versa.
+Unleashing the full potential of FAS necessitates the acquisition of $\mathbf{g}_{u}$.
+Conventional pilot-based approaches to acquire $\mathbf{g}_{u}$ cause excessive overhead [37].
+To this end, we propose channel extrapolation to acquire the CSI of a large set of ports $\widehat{\mathbf{g}}_{u,\mathcal{A}}$ using a small set of observable ports $\mathbf{g}_{u,\mathcal{B}}$, which is expressed as:
+$\widehat{\mathbf{g}}_{u,\mathcal{A}}=f(\mathbf{g}_{u,\mathcal{B}}),$ (3)
+where f()f(*) is the channel extrapolation function.
+The aim of the extrapolation is to minimize the expected normalized mean-square-error (NMSE) between the estimated CSI $\widehat{\mathbf{g}}_{u,\mathcal{A}}$ and the ground truth CSI ${\mathbf{g}}_{u,\mathcal{A}}$, which is expressed as:
+$\min_{f(*)}\mathbb{E}\left(\frac{|\widehat{\mathbf{g}}_{u,\mathcal{A}}-{\mathbf{g}}_{u,\mathcal{A}}|^{2}}{|{\mathbf{g}}_{u,\mathcal{A}}|^{2}}\right).$ (4)
+III Proposed Method
 Here, we propose CANet, a noise-robust architecture for the channel extrapolation task.
@@
 III-A Masking Strategy
-We denote the perfect CSI as $\mathbf{g}\in\mathbb{C}^{N_{s}\times M_{t}}$, and construct a tensor $\mathbf{u}\in\mathbb{R}^{2M_{t}\times N_{y}\times N_{x}}$ to represent the concatenated real and imaginary components of the CSI across a 2D fluid antenna grid of size $N_{x}\times N_{y}$.
+We denote the ground truth CSI as $\mathbf{g}\in\mathbb{C}^{N_{s}\times M_{t}}$, and construct a tensor $\mathbf{u}\in\mathbb{R}^{2M_{t}\times N_{y}\times N_{x}}$ to represent the concatenated real and imaginary components of the CSI across a 2D fluid antenna grid of size $N_{x}\times N_{y}$.
 To simulate partial observations, a subset of antenna ports is randomly masked along the spatial grid, with a masking ratio typically ranging from $80\%$ to $95\%$.
-The resulting masked tensor is denoted as $\tilde{\mathbf{u}}\in\mathbb{R}^{2M_{t}\times N_{y}\times N_{x}}$, in which the unobserved positions are replaced with a distinct negative value (e.g., $-10$) to make them distinguishable for the network.
-Additionally, we define a binary indicator matrix $\mathbf{E}_{\text{flag}}\in\{0,1\}^{N_{y}\times N_{x}}$, where a value of $1$ indicates an observed port and 00 is a masked (unobserved) port.
+The resulting masked tensor is denoted as $\tilde{\mathbf{u}}\in\mathbb{R}^{2M_{t}\times N_{y}\times N_{x}}$, in which the unobserved positions are replaced with a distinct negative value (e.g., 10-10) to make them distinguishable for the network.
+Additionally, we define a binary indicator matrix $\mathbf{E}_{\text{flag}}\in\{0,1\}^{N_{y}\times N_{x}}$, where a value of $1$ indicates an observed port and 0 is a masked (unobserved) port.
 No prior knowledge of the channel is used to select ports for extrapolation.
@@
 III-B Model Architecture
-The overall architecture of CANet consists of four training modules. Next, we summarize the role of each module.
+The architecture of CANet is designed to exploit the spatial continuity and spectral redundancy of FAS.
+It leverages a local-to-global inductive bias to reconstruct CSI from sparse observations, as detailed below.
 III-B1 CAB
-For channel extrapolation tasks with strong spatial correlations, using convolutional neural network is a natural choice for feature extraction.
-However, conventional convolution is susceptible to interference from unobserved ports within the receptive field, which can degrade the quality of feature representations.
-To overcome this issue, we propose a dual-branch architecture that separately models and predicts the CSI of observed and unobserved ports, thereby enhancing both the modeling accuracy and the robustness of extrapolation.
+To tackle the challenge of irregular sparse observations inherent in FAS, we design the CAB to precisely extract effective information within the receptive field.
+Standard convolution treats all pixels equally, making it difficult to distinguish valid signals from the values at unobserved positions.
+To overcome this, CAB employs a context-aware mechanism that dynamically identifies and weighs valid features from observed ports, ensuring accurate local feature extraction.
 Specifically, we first compute a soft mask (confidence map) over the entire feature map using convolution.
@@
 $x^{a}=\sum_{i\in\mathcal{B}_{o}}\left\langle mw^{q}x_{o},mw^{k}x_{i}\right\rangle mx_{i}w^{v}x_{i},$ (6)
-[…]
-$m\in\sum_{i\in\mathcal{B}_{o}}w_{i}^{c}x_{i}.$ (5)
+where $\left\langle\cdot,\cdot\right\rangle$ denotes the composite operation of dot-product with softmax.
+This self-attention mechanism empowers the model to focus on context information pertinent to the current pixel, thereby enhancing the capability of features to capture and discriminate CSI.
+In parallel, within the convolutional branch, we employ another set of convolution kernels $w^{c}$ to extract local channel features $x^{c}$ by
+$x^{c}=\sum_{i\in\mathcal{B}_{o}}w_{i}^{c}x_{i}.$ (7)
 Finally, we perform a weighted fusion of the convolutional branch output $x_{c}$ and the attention branch output $x_{a}$ based on the soft mask $m$, yielding the output
@@
 III-B2 CSCA
-In the CSCA module, we first reduce the input feature channels to half of the original before feeding them into two self-attention modules.
+Considering the finite aperture of FAS, multipath propagation creates specific geometric constraints and long-range correlations across the antenna surface.
+We introduce CSCA to capture these global dependencies that local convolutions might miss.
+By facilitating cross-scale contextual attention, it promotes global consistency of the extrapolated channel with the underlying propagation geometry over the finite aperture.
+We first reduce the input feature channels to half of the original before feeding them into two self-attention modules.
 One branch directly undergoes self-attention at the same scale, while the other duplicates the feature map and down-samples it to half of the original resolution.
@@
 At each self-attention stage, patches are extracted from the input features, and the cosine similarity between patches in the missing region and those in the external region is computed to establish cross-region feature correlations.
-Formally, let $q_{i}$ denote the $i$-th patch from the observed region $\psi^{l}$, and $k_{j}$ denote the $j$-th patch from the masked region $\phi^{l}$.
+Formally, let $q_{i}$ denote the $i$-th patch from the masked region $\psi^{l}$, and $k_{j}$ denote the $j$-th patch from the observed region $\phi^{l}$.
 The cosine similarity is computed as
@@
 where $v_{j}$ represents the $j$-th patch extracted from features outside unobserved regions. Finally we concatenate the results of the two branches as the final result.
-TABLE I: Architecture of the proposed CANet
-Layer Operator Input Size Output Size Output Channel Kernel Size Stride Nonlinearity
-1 CAB $32~{}\times~{}16$ $32~{}\times~{}16$ $64$ $5$ $1$ LeakyReLU
-2 CAB $32~{}\times~{}16$ $32~{}\times~{}16$ $128$ $3$ $2$ LeakyReLU
-3 CAB $32~{}\times~{}16$ $16~{}\times~{}8$ $128$ $3$ $1$ LeakyReLU
-4 CAB $16~{}\times~{}8$ $16~{}\times~{}8$ $256$ $3$ $1$ LeakyReLU
-5 CAB $16~{}\times~{}8$ $16~{}\times~{}8$ $256$ $3$ $1$ LeakyReLU
-6 CSCA $16~{}\times~{}8$ $16~{}\times~{}8$ $256$ Null Null Null
-7–16 ConvNeXt v2 $16~{}\times~{}8$ $16~{}\times~{}8$ $256$ $7$ $3$ GeLU
-17 BilinerUp $16~{}\times~{}8$ $32~{}\times~{}16$ $256$ $3$ $1$ Null
-18 CAB $32~{}\times~{}16$ $32~{}\times~{}16$ $128$ $3$ $1$ LeakyReLU
-19 CAB $32~{}\times~{}16$ $32~{}\times~{}16$ $128$ $3$ $1$ LeakyReLU
-20 CAB $32~{}\times~{}16$ $32~{}\times~{}16$ $64$ $3$ $1$ LeakyReLU
-21 CAB $32~{}\times~{}16$ $32~{}\times~{}16$ $64$ $3$ $1$ LeakyReLU
-22 Conv $32~{}\times~{}16$ $32~{}\times~{}16$ $16$ $7$ $1$ Null
 III-B3 ConvNeXt v2 with dropout regularization
-After CAB and CSCA, we adopt ConvNeXt V2 [29] with dropout regularization as the backbone network to further enhance feature extraction and improve generalization.
-ConvNeXt has been widely used for its strong feature representation capabilities, and ConvNeXt V2 builds upon this by introducing the global response normalization (GRN) module, which enables more expressive feature representations.
-In our design, we employ the ConvNeXt V2 block for channel extrapolation.
+After the initial completion by CAB and CSCA, we employ the ConvNeXt v2 [31] backbone to perform deep feature refinement and learn robust high-level channel structures.
+ConvNeXt is known for its strong representation capabilities, and ConvNeXt v2 further enhances this by introducing the global response normalization (GRN) module, which enables more expressive feature learning.
+In our design, we employ the ConvNeXt v2 block integrated with dropout regularization to prevent overfitting and ensure generalization under diverse noise conditions.
 By randomly dropping a subset of neurons during training, dropout alters the effective network structure in each training epoch, leading to a degree of statistical independence across different training instances, thereby reducing prediction variance and enhancing generalization under diverse noise conditions.
 III-B4 Spatial amplitude perturbation
+Since FAS channel extrapolation is highly sensitive to measurement noise and multipath variations, we introduce a Spatial Amplitude Perturbation module to enforce robustness by simulating spectral uncertainty.
 To enhance the model’s robustness, we propose a spatial amplitude perturbation strategy at the input stage, inspired by frequency perturbation techniques in the image domain [18, 35].
-Specifically, the input CSI features are first zero-mean normalized to remove the DC component, then transformed into the frequency domain via FFT.
+Specifically, the input CSI features are first zero-mean normalized to remove the direct current (DC) component, then transformed into the frequency domain via FFT.
 The amplitude spectrum is then perturbed in a stochastic manner based on a predefined probability, effectively simulating potential frequency interference and noise.
@@
 $\hat{\mathbf{z}}=\mathcal{F}^{-1}\left(\mathcal{F}(\tilde{\mathbf{z}})\cdot\left(1+\bm{\gamma}\cdot\bm{\xi}\cdot\bm{\delta}\right)\right)+\bar{\mathbf{z}},$ (13)
-where $\tilde{\mathbf{z}}$ is the spatially zero-mean feature, $\hat{\mathbf{z}}$ is the final perturbed feature after reconstruction, $\mathcal{F}(\cdot)$ and $\mathcal{F}^{-1}(\cdot)$ denote the 2D discrete Fourier transform and its inverse.
+where $\tilde{\mathbf{z}}$ is the spatially zero-mean feature, $\hat{\mathbf{z}}$ is the final perturbed feature after reconstruction, $\bar{\mathbf{z}}$ denotes the spatial mean component removed in (12), $\mathcal{F}(\cdot)$ and $\mathcal{F}^{-1}(\cdot)$ denote the 2D discrete Fourier transform and its inverse.
 The perturbation mask $\bm{\delta}\in\{0,1\}^{H\times W}$ above is sampled from a Bernoulli distribution with probability $\mu$, and the amplitude coefficients $\bm{\xi}\in[0,1]^{H\times W}$ are sampled from a normal distribution $\mathcal{N}(0,1)$.
@@
 $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{MSE}}+\beta\mathcal{L}_{\mathrm{FFT}},$ (14)
-where $\mathcal{L}_{\text{MSE}}$ and $\mathcal{L}_{\mathrm{FFT}}$ are the mean-square-error (MSE) loss and the FFT loss attributed to spatial amplitude perturbation, respectively, calculated as
+where $\mathcal{L}_{\text{MSE}}$ and $\mathcal{L}_{\mathrm{FFT}}$ are the mean squared error (MSE) loss and the FFT loss attributed to spatial amplitude perturbation, respectively, calculated as
 $\mathcal{L}_{\text{MSE}}$ $=\mathbb{E}\left[\frac{1}{|\Omega|}\sum_{(y,x)\in\Omega}\left\|\hat{\mathbf{{U}_{c}}}[:,y,x]-\mathbf{{U}_{c}}[:,y,x]\right\|_{2}^{2}\right],$ (15)
@@
 IV Numerical Results
+IV-A Simulation settings
+TABLE I: Architecture of the proposed CANet.
+Layers Operator Out Ch. K(S) Act.
+1 CAB $32{\times}16$ 64 $5(1)$ LReLU
+2 CAB $32{\times}16$ 128 $3(2)$ LReLU
+3 CAB $16{\times}8$ 128 $3(1)$ LReLU
+4–5 CAB $\times 2$ $16{\times}8$ 256 $3(1)$ LReLU
+6 CSCA $16{\times}8$ 256
+7–16 ConvNeXt v2 $\times 10$ $16{\times}8$ 256 $7(3)$ GELU
+17 BilinearUp $32{\times}16$ 256
+18–19 CAB $\times 2$ $32{\times}16$ 128 $3(1)$ LReLU
+20–21 CAB $\times 2$ $32{\times}16$ 64 $3(1)$ LReLU
+22 Conv $32{\times}16$ 16 $7(1)$
+Note: K: kernel size; S: stride; LReLU: LeakyReLU; GELU: Gaussian Error Linear Unit.
 In the simulations, the FAS channel involves $U_{t}$ users.
@@
 The total number of ports at the FAS side is set to $N_{s}=512$.
-The performance is evaluated through the normalized MSE (NMSE), defined as
+The performance is evaluated through the NMSE, defined as
 $\text{NMSE}(\hat{\mathbf{g}},\mathbf{g})=\frac{\sum_{n=1}^{N_{\text{test}}}\left\|\mathbf{g}_{n}-\hat{\mathbf{g}}_{n}\right\|^{2}}{\sum_{n=1}^{N_{\text{test}}}\left\|\mathbf{g}_{n}\right\|^{2}},$ (17)
@@
 We set the maximum strength $\gamma$ to $0.5$, $\mu$ to $0.05$, and the dropout rate to $0.5$, $\beta$ to $0.02$.
-(a) $W_{s}=2~{}{\rm cm}\times 4~{}{\rm cm}$
-(b) $W_{s}=8~{}{\rm cm}\times 16~{}{\rm cm}$
-Figure 2: NMSE versus the number of observed ports with different FAS sizes.
+(a) 3.4GHz
+(b) 28GHz
+Figure 2: NMSE versus the number of observed ports with different operating frequencies.
+CANet consistently outperforms benchmark models across all observation ratios, demonstrating superior robustness in reconstructing high-resolution CSI from sparse observations.
+IV-B Performance analysis
+Fig.
+2 illustrates the NMSE performance of different models as a function of the observation ratio for a $2~\mathrm{cm}\times 4~\mathrm{cm}$ FAS array under multiple SNR conditions, where Figs.
+2(a) and 2(b) correspond to the carrier frequencies of $3.4$ GHz and $28$ GHz, respectively.
+We reproduce AGMAE in [37] as a representative deep-learning-based baseline tailored for FAS scenarios, and additionally include CT[39] as a comparison model, which was originally designed for channel extrapolation in conventional multiple-input multiple-output (MIMO) systems, to evaluate the applicability of generic extrapolation models to FAS settings.
+In the figure, the same line style is used to indicate the same SNR condition.
+For both carrier frequencies, the NMSE of all models decreases monotonically with the observation ratio, indicating that richer port observations improve CSI reconstruction accuracy.
+However, under sparse observations, especially in low-SNR conditions, CT exhibits limited performance gains, suggesting that directly reusing MIMO-oriented extrapolation structures is insufficient to capture the fine-grained port correlations inherent in FAS.
+In contrast, AGMAE demonstrates more stable extrapolation performance across both frequency bands, while the proposed CANet consistently achieves the lowest NMSE under all SNR and observation ratio settings, highlighting its adaptability to different carrier frequencies.
+To further evaluate the impact of channel extrapolation accuracy on communication performance, the outage probability is evaluated based on the signal model and port selection criteria described in [27].
+Fig.
+3 presents the outage probability (OP) as a function of the SINR threshold under a fixed observation ratio of $10\%$ for different methods.
+Here, Baseline denotes the case where transmission is performed using full but noisy CSI, while Reference corresponds to direct transmission based only on the observable port CSI.
 Figs.
-2(a) and 2(b) illustrate the NMSE curves of different models under various SNR conditions, as a function of the observed port ratio, for FAS sizes of $2~{}{\rm cm}\times 4~{}{\rm cm}$ and $8~{}{\rm cm}\times 16~{}{\rm cm}$, respectively.
-We reproduce the AGMAE in [25] as a baseline for comparison.
-The simplified version of CANet without spatial amplitude perturbation and spectral consistency loss is denoted as ‘CANet-B’, while the full CANet model includes all the proposed components.
-In the figures, curves with the same line style represent the same SNR setting.
-The results demonstrate that as the ratio of observed ports increases, the NMSE of all the models decreases considerably, showing that richer observed information leads to better reconstruction performance.
-Under extremely low observation rates (e.g., $5\%$), AGMAE slightly outperforms CANet-B, suggesting that its structure has certain advantages under highly sparse inputs.
-But in most observation rates, CANet-B consistently surpasses AGMAE.
-Furthermore, with the incorporation of amplitude spectrum perturbation and spectral consistency loss, the overall performance of CANet is further improved.
-For example, at an observation rate of $20\%$ with the $2~{}{\rm cm}\times 4~{}{\rm cm}$ FAS, CANet achieves up to a $3~{}{\rm dB}$ NMSE gain over CANet-B, demonstrating stronger noise robustness and generalization ability.
-Fig.
-2(b) shows a similar trend under a larger FAS size, validating the adaptability and effectiveness of the proposed method across different physical FAS configurations.
+3 correspond to observation input SNRs of 0 dB and $20$ dB, respectively.
+Figure 3: Outage probability versus SINR threshold at a fixed observation ratio of $10\%$ under 0 dB and 20 dB observation SNR.
+CANet achieves the lowest outage probability among extrapolation schemes, effectively mitigating noise to approach the performance of direct transmission.
+Under both SNR settings, CANet achieves the lowest OP over the entire SINR threshold range.
+Although the Baseline exploits full-port observations, its performance is still limited by measurement noise; in contrast, CANet effectively suppresses observation noise during extrapolation by learning the spatial correlation and statistical structure of FAS channels, thereby achieving improved communication reliability.
+As the observation input SNR increases to $20$ dB, the performance gap among different methods narrows, while CANet consistently maintains a clear advantage, demonstrating its robustness under varying noise conditions.
+TABLE II: Ablation Study of Physics-Informed Modules ($2~\mathrm{cm}\times 4~\mathrm{cm}$ FAS, 20% Observation Ratio)
+ID Modules NMSE (dB)
+Perturb. FFT Loss 0 dB 10 dB 20 dB
+1 6.22-6.22 15.88-15.88 18.86-18.86
+2 6.51-6.51 16.42-16.42 19.45-19.45
+3 7.35-7.35 17.46-17.46 20.47-20.47
+4 $\textbf{8.50}$ $\textbf{18.85}$ $\textbf{21.83}$
+[…]
+IV-C Ablation experiments
+To verify the effectiveness of the proposed FAS-motivated components, an ablation study is conducted as shown in Table II.
+Compared to the basic model (ID 1), the independent inclusion of the FFT loss (ID 2) and the spatial amplitude perturbation strategy (ID 3) both lead to significant NMSE reductions across various SNR levels.
+Specifically, the full CANet (ID 4) achieves the best performance, with an additional gain of approximately 3 dB at 20 dB SNR compared to ID 1.
+These results confirm that explicitly enforcing frequency-domain consistency and simulating spatial uncertainty effectively guide the network to capture the intrinsic propagation physics rather than overfitting to pixel-wise noise.
 V Conclusions